A force of 50N is required to push a car on a level road with a constant speed of 10m/s. The mass of the car is 500kg. What force should be applied in the car to make it accelerate at 1m/s^2?

Grade:11

1 Answers

Adarsh
7 years ago
 
From the given data we will get the frictional force on the car while it is moving is 50N.
When we apply a force of magnitude 50N on it when it is moving with 10m/s in the direction of its motion, both the frictional force and the applied force gets cancelled and then the body will be moving with constant speed(10m/s).
Again, if we want to accelerate it with an acceleration 1m/s2, we have to apply an extra force of magnitude =ma = 500×1 = 500N in the direction of its motion.
So net force we have to apply = 50N + 500N = 550N.
 
Hope you understand my solution.
 
ALL THE BEST...